OpenStudy (anonymous):

Evaluate the expression: 2a + 4b – 3c for a = 9, b = 5, c = 12

OpenStudy (anonymous):

=2(9)+4(5)-3(12) =18+20-36 =38-36 =2

OpenStudy (anonymous):

Thank you. So much. That really helped me picture it, and put the pieces together.

OpenStudy (anonymous):

your welcome!!!